An opinion poll of registered voters is performed to estimate the support for an Important Ballot Initiative. Each person polled can indicate the y support the intialtive (S) or donot support it (D).

1.if N people are polled,write down a probability distribution function that should describe the chances of finding that n respondents support the initiative, if any given voter has unknown chance p of supporting the initiative.

2.The poll takes place. It is found that 88 people support and 95 people donot support. What is p as defined in PartA? What is the margin of error(in%)for the poll?
Answer #1

1) Probability function would be =

2) N = 88+95 =

n = 88

p = 88/183 = 0.4809

Margin of error =

= 7.239%